A POPULATION CONSISTS OF THE NUMBERS 2, 4, 9, 10 AND 5. LET US LIST ALL POSSIBLE SAMPLE SIZE OF 3 FROM THIS

POPULATION AND COMPUTE THE MEAN OF EACH SAMPLE. COMPUTE FOR THE POPUALATION MEAN.


1
Expert's answer
2021-11-29T19:06:15-0500

Sample Mean

2,4,9 5

2,4,10 16/3

2,4,5 11/3

2,9,10 7

2,9,5 16/3

2,10,5 17/3

4,9,10 23/3

4,9,5 6

4,10,5 19/3

9,10,5 8


Population mean: "\mu=\frac{2+4+9+10+5}{5}=6."

Mean of the sample distribution:

"\mu_{\bar x}=\frac{1}{10}(5+\frac{16}{3}+\frac{11}{3}+7+\frac{16}{3}+\frac{17}{3}+\frac{23}{3}+6+\frac{11}{3}+8)=6."